# Asymmetric Errors

### v1

@@ B.6 The matched quintic
The matched quintic log likelihood model assumes that, outside of the interval $[-{\sigma^{-}},{\sigma^{+}}]$, the log likelihood behaves as a second order polynomial with the second derivative $-1/{\sigma^{+}}^{2}$ for $a>{\sigma^{+}}$ and $-1/{\sigma^{-}}^{2}$ for $a<-{\sigma^{-}}$. Inside the interval $[-{\sigma^{-}},{\sigma^{+}}]$ the polynomial takes the form
@@
$\begin{split}\alpha&=-10({\sigma^{-}}-{\sigma^{+}})/\eta\\ \beta&=-18({\sigma^{-}}-{\sigma^{+}})^{2}/\eta\\ \gamma&=45{\sigma^{-}}{\sigma^{+}}({\sigma^{-}}-{\sigma^{+}})/\eta\\ \delta&=(8{\sigma^{-}}^{4}+19{\sigma^{-}}^{3}{\sigma^{+}}-19{\sigma^{-}}^{2}{\sigma^{+}}^{2}+19{\sigma^{-}}{\sigma^{+}}^{3}+8{\sigma^{+}}^{4})/\eta\end{split}$ (64)
@@
This log likelihood model behaves similarly to Eq. 57 for large values of $|a|$ but avoids the discontinuity in the second derivative (the third derivative is still discontinuous at $a=-{\sigma^{-}}$ and $a={\sigma^{+}}$).
@@
In combination with Eq. 57 for $a<-{\sigma^{-}}$ and $a>{\sigma^{+}}$, this model has a very reasonable behavior for large values of $|a|$ but can be somewhat ”wavy” on the interval $[-{\sigma^{-}},{\sigma^{+}}]$. It also has a discontinuous third derivative at $a=-{\sigma^{-}}$ and $a={\sigma^{+}}$.
@@ B.8 The double quartic
$\ln L(a)=\begin{cases}P_{2}^{-}(a)&\text{for }a\leq-{\sigma^{-}}\\ P_{4}^{-}(a)&\text{for }-{\sigma^{-}}\leq a\leq 0\\ P_{4}^{+}(a)&\text{for }0\leq a\leq{\sigma^{+}}\\ P_{2}^{+}(a)&\text{for }{\sigma^{+}}\leq a\end{cases}$ (67)
where $P_{2}^{-}(a)$ and $P_{2}^{+}(a)$ are quadratic polynomials and $P_{4}^{-}(a)$ and $P_{4}^{+}(a)$ are quartics defined by the following conditions: $P_{2}^{-}(-{\sigma^{-}})=P_{4}^{-}(-{\sigma^{-}})=P_{2}^{+}({\sigma^{+}})=P_{4}^{+}({\sigma^{+}})=-\frac{1}{2}$, $\left.\frac{d}{da}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d}{da}P_{4}^{-}(a)\right|_{a=-{\sigma^{-}}}$, $\left.\frac{d^{2}}{da^{2}}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d^{2}}{da^{2}}P_{4}^{-}(a)\right|_{a=-{\sigma^{-}}}=-\frac{1}{{\sigma^{-}}^{2}}$, $\left.\frac{d}{da}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d}{da}P_{4}^{+}(a)\right|_{a={\sigma^{+}}}$, $\left.\frac{d^{2}}{da^{2}}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d^{2}}{da^{2}}P_{4}^{+}(a)\right|_{a={\sigma^{+}}}=-\frac{1}{{\sigma^{+}}^{2}}$, $P_{4}^{-}(0)=P_{4}^{+}(0)=0$, $\left.\frac{d}{da}P_{4}^{-}(0)\right|_{a=0}=\left.\frac{d}{da}P_{4}^{+}(0)\right|_{a=0}=0$, $\left.\frac{d^{2}}{da^{2}}P_{4}^{-}(0)\right|_{a=0}=\left.\frac{d^{2}}{da^{2}}P_{4}^{+}(0)\right|_{a=0}=-\frac{1}{\sigma_{0}^{2}}$. When the value of the parameter $\sigma_{0}$ is selected, these conditions define the polynomial coefficients uniquely. The resulting curve is continuous together with its first two derivatives.
Two choices of $\sigma_{0}$ have been investigated. The simple double quartic is obtained by setting $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$. For this choice of $\sigma_{0}$, the requirement that $\ln L(a)$ has a single extremum (the maximum at 0) leads to the condition that ${\sigma^{+}}$ and ${\sigma^{-}}$ must differ by less than a factor of $\frac{68}{11}$.
@@ B.9 The double quintic
where $P_{2}^{-}(a)$ and $P_{2}^{+}(a)$ are quadratic polynomials and $P_{5}^{-}(a)$ and $P_{5}^{+}(a)$ are quintics defined by the following conditions: $P_{2}^{-}(-{\sigma^{-}})=P_{5}^{-}(-{\sigma^{-}})=P_{2}^{+}({\sigma^{+}})=P_{5}^{+}({\sigma^{+}})=-\frac{1}{2}$, $\left.\frac{d}{da}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d}{da}P_{5}^{-}(a)\right|_{a=-{\sigma^{-}}}=\frac{1}{{\sigma^{-}}}$, $\left.\frac{d^{2}}{da^{2}}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d^{2}}{da^{2}}P_{5}^{-}(a)\right|_{a=-{\sigma^{-}}}=-\frac{1}{{\sigma^{-}}^{2}}$, $\left.\frac{d}{da}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d}{da}P_{5}^{+}(a)\right|_{a={\sigma^{+}}}=-\frac{1}{{\sigma^{+}}}$, $\left.\frac{d^{2}}{da^{2}}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d^{2}}{da^{2}}P_{5}^{+}(a)\right|_{a={\sigma^{+}}}=-\frac{1}{{\sigma^{+}}^{2}}$, $P_{5}^{-}(0)=P_{5}^{+}(0)=0$, $\left.\frac{d}{da}P_{5}^{-}(0)\right|_{a=0}=\left.\frac{d}{da}P_{5}^{+}(0)\right|_{a=0}=0$, $\left.\frac{d^{2}}{da^{2}}P_{5}^{-}(0)\right|_{a=0}=\left.\frac{d^{2}}{da^{2}}P_{5}^{+}(0)\right|_{a=0}=-\frac{1}{\sigma_{0}^{2}}$. When the value of the parameter $\sigma_{0}$ is selected, these conditions define the polynomial coefficients uniquely. The resulting curve is continuous together with its first two derivatives. It coincides with Eq. 57 outside of the interval $[-{\sigma^{-}},{\sigma^{+}}]$.
Two choices of $\sigma_{0}$ have been investigated. The simple double quintic is obtained by setting $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$. For this choice of $\sigma_{0}$, the requirement that $\ln L(a)$ has a single extremum (the maximum at 0) leads to the condition that ${\sigma^{+}}$ and ${\sigma^{-}}$ must differ by less than a factor of $13.5$.
@@ B.10 The conservative spline
$\ln L(a)=\begin{cases}P_{2}^{-}(a)&\text{for }a\leq a_{\mbox{\scriptsize\,left}}\\ \alpha a^{3}+\beta a^{2}&\text{for }a_{\mbox{\scriptsize\,left}}\leq a\leq a_{\mbox{\scriptsize\,right}}\\ P_{2}^{+}(a)&\text{for }a\geq a_{\mbox{\scriptsize\,right}}\end{cases}$ (69)
@@
This log-likelihood construction mitigates the discontinuity of the second derivative afflicting Eq. 57 while maintaining user-controlled lower and upper bounds on the Fisher information of the log-likelihood in the parameter space. If a linear log-likelihood curve is added to this one, the uncertainties of the result determined from the $\Delta\ln L(a)=-\frac{1}{2}$ equation will be bounded by the range $[\mbox{min}({\sigma^{-}},{\sigma^{+}})/\sqrt{\kappa},\ \sqrt{\kappa}\,\mbox{max}({\sigma^{-}},{\sigma^{+}})]$ no matter how large is the slope of the added curve.
@@ B.11 The log logistic-beta
The density of the logistic-beta distribution (a.k.a. the type IV generalized logistic) is given by [ref:johnsonkotzv2]:
@@
$\ln L(a)=gc^{2}\left[(A-1)\ln\left(\frac{1}{2}\left(A\left(e^{\frac{a}{c}}-1\right)+e^{\frac{a}{c}}+1\right)\right)-(A+1)\ln\left(\frac{1}{2}e^{-\frac{a}{c}}\left(A\left(e^{\frac{a}{c}}-1\right)+e^{\frac{a}{c}}+1\right)\right)\right].$ (71)
@@
  • $g$ is the parameter regulating the magnitude of $\frac{d^{2}\ln L(a)}{da^{2}}$ maximum. Indeed, it can be shown that $\ln L(a)$ is strictly concave, while $\left|\frac{d^{2}\ln L(a)}{da^{2}}\right|$ is maximized at $a=c\ln\left(\frac{1-A}{1+A}\right)$. At that point, $\frac{d^{2}\ln L(a)}{da^{2}}=-\frac{g}{2}$.
@@
Eq. 71 satisfies the conditions $\ln L(0)=0$ and $\left.\frac{d\ln L(a)}{da}\right|_{a=0}=0$ automatically. It also needs to satisfy the conditions $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})=-\frac{1}{2}$. While with three parameters $A$, $c$, and $g$ the latter conditions can be satisfied in a multitude of ways, we choose the parameterization that minimizes $g$ (i.e., $A$ and $c$ are chosen in such a way that $|\ln L({\sigma^{+}})/g|$ is maximized subject to the constraint $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})$). This results in a conservative log-likelihood model with limited slopes and the second derivative exceeding $\frac{1}{\min({\sigma^{-}},{\sigma^{+}})^{2}}$ in magnitude only on a short interval.
@@ B.12 The logarithmic
$\beta={{\sigma^{+}}\over{\sigma^{-}}}\qquad\gamma={{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}{\sigma^{-}}}.$ (73)
@@ B.13 The generalised Poisson
$\ln L(a)=-\alpha(a+\beta)+{\cal N}\ln(a+\beta)+constant$ (74)
@@
$\ln L=-\alpha x+{\cal N}\left(1+{\alpha a\over{\cal N}}\right)$ (75)
@@
${1-\gamma{\sigma^{-}}\over 1+\gamma{\sigma^{+}}}=e^{-\gamma({\sigma^{+}}+{\sigma^{-}})}$ (76)
This has to be solved numerically for $\gamma$. The solution lies between 0 and $1/{\sigma^{-}}$, and can be found by repeatedly taking the midpoint. (Attempts to use more sophisticated algorithms were unsuccessful.) ${\cal N}$ is then found from
@@ B.14 The linear sigma
Bartlett suggests (and justifies) that the likelihood function is described by a Gaussian whose width varies as a function of the parameter [bartlett1, bartlett2, deltalnL]
@@
This does not include the $-\ln\sigma$ term from the denominator of the Gaussian, however omitting this term improves the accuracy of $\Delta\ln L=-{1\over 2}$ errors[deltalnL], bringing them into line with the Bartlett form.
If we suppose that the variation of $\sigma$ is linear, at any rate in the region of interest.
$\sigma(a)=\sigma+\sigma^{\prime}(a-\hat{a})$ (79)
$\ln L(a)=-{1\over 2}\left({a-\hat{a}\over\sigma+\sigma^{\prime}(a-\hat{a})}\right)^{2}$ (80)
@@
$\sigma={2{\sigma^{+}}{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}\qquad\sigma^{\prime}={{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}$ (81)

### v2

@@ B.6 The matched quintic
The matched quintic log likelihood model assumes that, outside of the interval $[-{\sigma^{-}},{\sigma^{+}}]$, the log likelihood behaves as a second order polynomial with the second derivative $-(1/{\sigma^{+}})^{2}$ for $a>{\sigma^{+}}$ and $-(1/{\sigma^{-}})^{2}$ for $a<-{\sigma^{-}}$. Inside the interval $[-{\sigma^{-}},{\sigma^{+}}]$ the polynomial takes the form
@@
$\begin{split}\alpha&=-10({\sigma^{-}}-{\sigma^{+}})/\eta\\ \beta&=-18({\sigma^{-}}-{\sigma^{+}})^{2}/\eta\\ \gamma&=45{\sigma^{-}}{\sigma^{+}}({\sigma^{-}}-{\sigma^{+}})/\eta\\ \delta&=(8{\sigma^{-}}^{4}+19{\sigma^{-}}^{3}{\sigma^{+}}-19{\sigma^{-}}^{2}{\sigma^{+}}^{2}+19{\sigma^{-}}{\sigma^{+}}^{3}+8{\sigma^{+}}^{4})/\eta\end{split}.$ (62)
@@
This log likelihood model behaves similarly to Equation 55 for large values of $|a|$ but avoids the discontinuity in the second derivative. The third derivative is still discontinuous at $a=-{\sigma^{-}}$ and $a={\sigma^{+}}$.
@@ B.7 The interpolated 7th degree polynomial
In combination with Equation 55 for $a<-{\sigma^{-}}$ and $a>{\sigma^{+}}$, this model has a very reasonable behavior for large values of $|a|$ but can be somewhat “wavy” on the interval $[-{\sigma^{-}},{\sigma^{+}}]$. It also has a discontinuous third derivative at $a=-{\sigma^{-}}$ and $a={\sigma^{+}}$.
@@ B.8 The double quartic
$\ln L(a)=\begin{cases}P_{2}^{-}(a)&\text{for }a\leq-{\sigma^{-}}\\ P_{4}^{-}(a)&\text{for }-{\sigma^{-}}\leq a\leq 0\\ P_{4}^{+}(a)&\text{for }0\leq a\leq{\sigma^{+}}\\ P_{2}^{+}(a)&\text{for }{\sigma^{+}}\leq a\end{cases},$ (65)
where $P_{2}^{-}(a)$ and $P_{2}^{+}(a)$ are quadratic polynomials and $P_{4}^{-}(a)$ and $P_{4}^{+}(a)$ are quartics defined by the following conditions:
$P_{2}^{-}(-{\sigma^{-}})=P_{4}^{-}(-{\sigma^{-}})=P_{2}^{+}({\sigma^{+}})=P_{4}^{+}({\sigma^{+}})=-\frac{1}{2},$
$\left.\frac{d}{da}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d}{da}P_{4}^{-}(a)\right|_{a=-{\sigma^{-}}},$
$\left.\frac{d^{2}}{da^{2}}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d^{2}}{da^{2}}P_{4}^{-}(a)\right|_{a=-{\sigma^{-}}}=-\frac{1}{{\sigma^{-}}^{2}},$
$\left.\frac{d}{da}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d}{da}P_{4}^{+}(a)\right|_{a={\sigma^{+}}},$ (66)
$\left.\frac{d^{2}}{da^{2}}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d^{2}}{da^{2}}P_{4}^{+}(a)\right|_{a={\sigma^{+}}}=-\frac{1}{{\sigma^{+}}^{2}},$
$P_{4}^{-}(0)=P_{4}^{+}(0)=0$ (67)
$\left.\frac{d}{da}P_{4}^{-}(0)\right|_{a=0}=\left.\frac{d}{da}P_{4}^{+}(0)\right|_{a=0}=0,$
$\left.\frac{d^{2}}{da^{2}}P_{4}^{-}(0)\right|_{a=0}=\left.\frac{d^{2}}{da^{2}}P_{4}^{+}(0)\right|_{a=0}=-\frac{1}{\sigma_{0}^{2}}.$
$\sigma_{0}$ is a somewhat arbitrary parameter describing the second differential at the peak. When it is selected, these conditions define the polynomial coefficients uniquely. The resulting curve is continuous together with its first two derivatives. Two choices of $\sigma_{0}$ have been investigated.
The simple double quartic is obtained by setting $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$. For this choice of $\sigma_{0}$, the requirement that $\ln L(a)$ has a single extremum (the maximum at 0) leads to the condition that ${\sigma^{+}}$ and ${\sigma^{-}}$ must differ by less than a factor of $\frac{68}{11}$.
@@ B.9 The double quintic
where $P_{2}^{-}(a)$ and $P_{2}^{+}(a)$ are quadratic polynomials and $P_{5}^{-}(a)$ and $P_{5}^{+}(a)$ are quintics defined by the following conditions:
$P_{2}^{-}(-{\sigma^{-}})=P_{5}^{-}(-{\sigma^{-}})=P_{2}^{+}({\sigma^{+}})=P_{5}^{+}({\sigma^{+}})=-\frac{1}{2},$
$\left.\frac{d}{da}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d}{da}P_{5}^{-}(a)\right|_{a=-{\sigma^{-}}}=\frac{1}{{\sigma^{-}}},$
$\left.\frac{d^{2}}{da^{2}}P_{2}^{-}(a)\right|_{a=-{\sigma^{-}}}=\left.\frac{d^{2}}{da^{2}}P_{5}^{-}(a)\right|_{a=-{\sigma^{-}}}=-\frac{1}{{\sigma^{-}}^{2}},$
$\left.\frac{d}{da}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d}{da}P_{5}^{+}(a)\right|_{a={\sigma^{+}}}=-\frac{1}{{\sigma^{+}}},$
$\left.\frac{d^{2}}{da^{2}}P_{2}^{+}(a)\right|_{a={\sigma^{+}}}=\left.\frac{d^{2}}{da^{2}}P_{5}^{+}(a)\right|_{a={\sigma^{+}}}=-\frac{1}{{\sigma^{+}}^{2}},$ (69)
$P_{5}^{-}(0)=P_{5}^{+}(0)=0,$
$\left.\frac{d}{da}P_{5}^{-}(0)\right|_{a=0}=\left.\frac{d}{da}P_{5}^{+}(0)\right|_{a=0}=0,$
$\left.\frac{d^{2}}{da^{2}}P_{5}^{-}(0)\right|_{a=0}=\left.\frac{d^{2}}{da^{2}}P_{5}^{+}(0)\right|_{a=0}=-\frac{1}{\sigma_{0}^{2}}.$
The resulting curve is continuous together with its first two derivatives. It coincides with Equation 55 outside of the interval $[-{\sigma^{-}},{\sigma^{+}}]$.
Again, when the value of the parameter $\sigma_{0}$ is selected, these conditions define the polynomial coefficients uniquely. Two choices of $\sigma_{0}$ have been investigated.
The simple double quintic is obtained by setting $\sigma_{0}=\sqrt{{\sigma^{-}}{\sigma^{+}}}$. For this choice of $\sigma_{0}$, the requirement that $\ln L(a)$ has a single extremum (the maximum at 0) leads to the condition that ${\sigma^{+}}$ and ${\sigma^{-}}$ must differ by less than a factor of $13.5$.
@@ B.10 The conservative spline
$\ln L(a)=\begin{cases}P_{2}^{-}(a)&\text{for }a\leq a_{\mbox{\scriptsize\,left}}\\ \alpha a^{3}+\beta a^{2}&\text{for }a_{\mbox{\scriptsize\,left}}\leq a\leq a_{\mbox{\scriptsize\,right}}\\ P_{2}^{+}(a)&\text{for }a\geq a_{\mbox{\scriptsize\,right}}\end{cases},$ (70)
@@
This log-likelihood construction mitigates the discontinuity of the second derivative afflicting Equation 55 while maintaining user-controlled lower and upper bounds on the Fisher information of the log-likelihood in the parameter space. If a linear log-likelihood curve is added to this one, the uncertainties of the result determined from the $\Delta\ln L(a)=-\frac{1}{2}$ equation will be bounded by the range $[\mbox{min}({\sigma^{-}},{\sigma^{+}})/\sqrt{\kappa},\ \sqrt{\kappa}\,\mbox{max}({\sigma^{-}},{\sigma^{+}})]$ no matter how large is the slope of the added curve.
@@ B.11 The log logistic-beta
The density of the logistic-beta distribution (a.k.a. the type IV generalized logistic) is given by [27]:
@@
$\ln L(a)=gc^{2}\left((A-1)\ln\left[\frac{1}{2}\left\{A\left(e^{\frac{a}{c}}-1\right)+e^{\frac{a}{c}}+1\right\}\right]-(A+1)\ln\left[\frac{1}{2}e^{-\frac{a}{c}}\left\{A\left(e^{\frac{a}{c}}-1\right)+e^{\frac{a}{c}}+1\right\}\right]\right).$ (72)
@@
  • $g$ is the parameter regulating the magnitude of the $\frac{d^{2}\ln L(a)}{da^{2}}$ maximum. Indeed, it can be shown that $\ln L(a)$ is strictly concave, while $\left|\frac{d^{2}\ln L(a)}{da^{2}}\right|$ is maximized at $a=c\ln\left(\frac{1-A}{1+A}\right)$. At that point, $\frac{d^{2}\ln L(a)}{da^{2}}=-\frac{g}{2}$.
@@
Equation 72 satisfies the conditions $\ln L(0)=0$ and $\left.\frac{d\ln L(a)}{da}\right|_{a=0}=0$ automatically. It also needs to satisfy the conditions $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})=-\frac{1}{2}$. While with three parameters $A$, $c$, and $g$ the latter conditions can be satisfied in a multitude of ways, we choose the parameterization that minimizes $g$, i.e. $A$ and $c$ are chosen in such a way that $|\ln L({\sigma^{+}})/g|$ is maximized subject to the constraint $\ln L(-{\sigma^{-}})=\ln L({\sigma^{+}})$. This results in a conservative log-likelihood model with limited slopes and the second derivative exceeding $\frac{1}{\min({\sigma^{-}},{\sigma^{+}})^{2}}$ in magnitude only on a short interval.
@@ B.12 The logarithmic
$\beta={{\sigma^{+}}\over{\sigma^{-}}}\qquad{\rm and}\qquad\gamma={{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}{\sigma^{-}}}.$ (74)
@@ B.13 The generalised Poisson
$\ln L(a)=-\alpha(a+\beta)+{\cal N}\ln(a+\beta)+constant,$ (75)
@@
$\ln L=-\alpha a+{\cal N}\ln\left(1+{\alpha a\over{\cal N}}\right).$ (76)
@@
${1-\gamma{\sigma^{-}}\over 1+\gamma{\sigma^{+}}}=e^{-\gamma({\sigma^{+}}+{\sigma^{-}})}.$ (77)
This has to be solved numerically for $\gamma$. The solution lies between 0 and $1/{\sigma^{-}}$, and can be found by repeatedly taking the midpoint.  [^4] Attempts to use more sophisticated algorithms were unsuccessful. ${\cal N}$ is then found from
@@ B.14 The linear sigma
Bartlett suggests (and justifies) that the likelihood function is described by a Gaussian whose width varies as a function of the parameter [16, 17, 13]
@@
This does not include the $-\ln\sigma$ term from the denominator of the Gaussian, however omitting this term improves the accuracy of $\Delta\ln L=-{1\over 2}$ errors[13].
If we suppose that the variation of $\sigma$ is linear, at any rate in the region of interest
$\sigma(a)=\sigma+\sigma^{\prime}(a-\hat{a}),$ (80)
$\ln L(a)=-{1\over 2}\left({a-\hat{a}\over\sigma+\sigma^{\prime}(a-\hat{a})}\right)^{2},$ (81)
@@
$\sigma={2{\sigma^{+}}{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}\qquad{\rm and}\qquad\sigma^{\prime}={{\sigma^{+}}-{\sigma^{-}}\over{\sigma^{+}}+{\sigma^{-}}}.$ (82)